The linear function f(x)=340x +4562 provides a model for the number
of students attending a community college from 1995 to 2005. If x = 0
corresponds to 1995, use this model to find the number of students who
attended the college in 2001.

Answer by TimothyLamb(4379)

---
f(x) = 340x + 4562
where:
x = year - 1995
---
find f(x) for year = 2001:
x = 2001 - 1995
x = 6
f(6) = 340(6) + 4562
---
answer:
f(6) = 6602
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php